• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name silwal

    Log entry time 07:55:59 on October 6,2009

    Entry number 293418

    keyword=owl shift summary

    00:00 taking production data @ 100 uA
    1:16 left arm Q1 tripped on low He levels, beam trip on ion chambers
    1:21 paged tech on call, Jack, asked MCC not to send us beam until Q1 is back up
    1:30 reset Q1, it is ramping up fine, will take about 15-20 minutes to get back up all the way
    1:36 asked MCC for 100 uA beam
    1:40 Q1 is fully back up.
    2:05 MCC checking some orbits at the injector, couple of minutes before beam is back
    2:31 MCC is beam back into the hall
    2:35 97 uA beam back, sort of, beam trips still
    3:16 MCC is taking the beam away for injector studies. they gave us about 30 minutes for the beam to be back
    3:25 MCC is sending us beam back
    3:50 MCC crew chief called to say that they were done adjusting the beam in the injector. they were having problems supplying both hall A and C their nominal currents at the same time. so they steered beam at the injector to get around it.
    4:20 left dipole down, requested escorted access into the hall.
    5:00 escorted access to the hall, I could not reset the dipole following instructions on the dipole. paged Jack. he is on his way.
    5:30 asked MCC for the beam, problem with left dipole, Jack could not bring it back up, need more careful study, probably a longer access to access the situation.
    6:15 bcm calibration finished, back to production with 100 uA
    7:45 all magnets down, cryo trip. cryo still unstable for left Q1, right Q1 and right Dipole.

    run # 12984: production run, left Q1 tripped towards the end of the run
    run # 12985: left Q1 ramping up, production run otherwise
    run # 12986: production, lots of beam trips
    run # 12987: LED test run, for det linearity test, no beam.
    run # 12988: production, left D1 went down towards the end of the run
    run # 12989: pedestal
    run # 12990: pedestal, short run
    run # 12991: production, left D1 down still, short run
    run # 12992: bcm calibration run
    run # 12993: CODA hangup
    run # 12994: production @ 100 uA, left D1 down still